Hi Tutors:
I have been working on this system in three for a few days. Please show me to solve with elimination:
4x-3y+z=16
x+2y-z=-3
3x+y+2z=1
Thanks!

Answer by longjonsilver(2297)

4x-3y+z=16 -- eqn1
x+2y-z=-3 -- eqn2
3x+y+2z=1 -- eqn3

Add 1 and 2 to give 5x-y = 13 --eqn4
Multiply all terms in eqn1 and then subtract eqn3 from it to give 5x - 7y = 31 --eqn5.

Now, subtract 5 from 4 to give 6y = -18 --> y = -3

now sub this into eqn4 or 5 to give x=2.

Now sub these 2 into eqn 1, 2 or 3, to find z. z=-1.


Check all 3 original equations work.

jon.